OpenStudy (anonymous):

Input the equation of the given line. The line including (4, 7) and (0, 7).

OpenStudy (anonymous):

find the slope using slope formula : (y2 - y1) / (x2 - x1) (4,7) x1 = 4 and y1 = 7 (0,7) x2 = 0 and y2 = 7 now we sub slope(m) = (7-7) / (0-4) slope(m) = 0/-4 slope(m) = 0 horizontal lines have a zero slope. Your line is horizontal and the equation for that line is : y = 7

OpenStudy (anonymous):

How would I write the equation?

OpenStudy (anonymous):

Sorry, one of my other questions was up and i was looking at that one. thank you.

OpenStudy (anonymous):

no problem :)